Vinyl flooring costs $25 per square yard. How much will it cost to buy vinyl flooring for a kitchen that is 4.5 yards wide and 7.5 yards long?
Answer by checkley75(3666)

4.5*7.5=33.75*25=$843.75 COST OF VINYL FLOORING FOR THE 4.5 BY 7.5 YARD ROOM.